# Traffic Response Functions: Patterns, Propagation and Congestion

### v1

@@ 4.2 On- and off-ramp propagation of congestion and generic response features
Besides the results of Sec. 4.1 we are interested in the capability of response functions to detect the propagation of congestion through on- and off-ramps into other parts of the motorway network. Therefore we take a look at two cases at interchange Breitscheid and Herne. In both cases we calculate the responses of spatially subsequent sections along an on-ramp or off-ramp, respectively, to an indicator section on a congested motorway. For interchange Breitscheid we chose section 1 as the indicator section and sections 2 and 6-8 as responding sections, where section 7 is located on the on-ramp leading traffic from motorway A3 onto motorway A52 and section 8 located on motorway A3 (see Fig. 2(a)). In case of interchange Herne we select section 9 as the indicator section and sections 10, 12 and 13 as responding sections, where section 12 is located on the on-ramp leading traffic from motorway A43 onto A42 (see Fig. 2(b)). Figure 5 contains the resulting responses for both cases.
As expected, one initially finds identical responses of all types for section 2
@@
at interchange Breitscheid in Fig. 5(a) (compared with Fig. 4(a)). Furthermore one can identify the propagation of congestion through the on- or off-ramp, respectively, considering the responses of all three traffic variables at sections 6 and 7 to all indicators. Here, the indicating minima (maxima) for $\varepsilon^{(0)}(t)$ are the most distinct. The responses for section 8 do not show any local minima or maxima, implying the congestion does not propagate to section 8. It is worth mentioning that there is no available detector between sections 7 and 8.
@@
In case of interchange Herne one finds a clear indication of congestion propagating through the on- or off-ramp at section 12 regarding the velocity responses to indicator $\varepsilon^{(0)}(t)$ in Fig. 5(b). The comparison between the latter and the corresponding flow responses reveals the incapability of the flow data to identify the congestion propagation in this case. While the flow responses of sections 12 and 13 exhibit a very fluctuating course without any clear local extrema, the responses of section 10 exhibit local minima (and maxima in case of indicator $\varepsilon^{(2)}(t)$). Considering the corresponding density responses one finds an anomaly. The courses of the density responses of sections 12 and 13 to indicator $\varepsilon^{(0)}(t)$ exhibit local maxima fitting the courses of the corresponding velocity responses and therefore the propagation of congestion. In contrast to this, the density response at section 10 exhibits a local minimum instead of an expected local maximum.
@@
Regardless of the possibility to clearly identify congestion propagation through on- or off-ramps in both cases in Fig. 5, we find the unexpected result of the density response for section 10 at interchange Herne. Without any obvious reason from a traffic point of view, or systematic errors, the unexpected result may originate from missing values in the analyzed data. For the indicator section 9 about 14% of that data are missing. For section 10 only 4% of the flow data and 9% of the velocity data is missing. The very fluctuating courses of the flow responses also suggest the potential influence of a low data quality on the anomalous results. Unfortunately, we can not corroborate this point due to a lack of other available data, e.g. on construction sites.
@@
Apart from the congestion propagation, the responses in Fig. 5 and Fig. 4 show similar courses for large time lags $\tau$. Comparing all studied cases one finds velocity responses increasing while flow and density responses decrease with growing time lag $\tau$. Our results imply that this behavior is a generic feature for the individual response types induced by the indicators defined in Eq. (6). These generic courses can be explained by the usual traffic situation represented by the averaged data in Fig. 3. With the increase of time the traffic situation during the rush hours calms down gradually, i.e. the number of traffic participants decreases. As a consequence, a decrease of density and flow enables higher velocities for the remaining traffic participants until the end of rush hours. The generic feature we find is suitable for all responses in our study.
@@ 4.3 Differences in congestion patterns
In Sec. 4.1 we find deviant behavior of responses to indicator $\varepsilon^{(2)}(t)$ compared to the other indicators. The deviant behavior clearly indicates an acceleration phase of the vehicles to moderate velocities for sections at interchange Herne in Fig. 4(b) and motorway A3 in Fig. 4(c), but not in case of interchange Breitscheid in Fig. 4(a). The consideration of these results raises the question why this is the case. From a traffic point of view, the results reveal the presence of different congestion patterns. In case of motorway A3 or interchange Herne vehicles are able to accelerate to higher velocities in range of indicator $\varepsilon^{(2)}(t)$ before decelerating again due to the present congestion. As a consequence larger gaps between the traffic participants have to be given, or else vehicles would not be able to accelerate into a higher velocity range. Therefore, the extrema of indicators $\varepsilon^{(0)}(t)$ and $\varepsilon^{(2)}(t)$ that occur at the same time shifts reveal a stop-and-go nature of the congestion pattern at these locations, i.e acceleration-deceleration cycles. In contrast, the responses for indicators $\varepsilon^{(1)}(t)$ and $\varepsilon^{(2)}(t)$ at interchange Breitscheid do not show any sign of an acceleration phase, i.e. occurring local maxima, during the crucial time period of congestion. This fact indicates a far more dense or synchronized nature in terms of congestion pattern, i.e. vehicles moving with low velocities (in range of $\varepsilon^{(0)}(t)$ and $\varepsilon^{(1)}(t)$ ) while maintaining small distances to vehicles driving ahead.
@@
This does not necessarily exclude a stop-and-go pattern of congestion, but the effect is not strong enough to be captured by the responses.
Due to the nature of the data at hand, it is not possible to verify the above assumptions with information about the distances between traffic participants, therefore we consider the raw data itself. Figure 6 shows a sample of the daily data according to Eq. (4) for all studied locations. In all cases the time period of traffic breakdown can clearly be recognized by considering the course of velocity and density. Due to its very fluctuating nature, the impact on traffic flow is not as clearly visible as in the case of the other observables. The velocity data (first column in Fig. 6) reveals what the averaged responses already indicated. While the velocity at section 3 at interchange Breitscheid almost invariably lies below the range of indicator $\varepsilon^{(2)}(t)$, the velocities at sections 12 and 15 at interchange Herne and on motorway A3, respectively, reach (and exceed) the range of indicator $\varepsilon^{(2)}(t)$ frequently. Even though a traffic oscillation, i.e. a stop-and-go pattern, is clearly recognizable in the velocities at all three locations, the corresponding responses in Fig. 4 only show an acceleration phase, i.e. local maxima, for indicator $\varepsilon^{(2)}(t)$ in case of interchange Herne and motorway A3. Considering the velocity at section 3 at Breitscheid in Fig. 6 one may expect to see a local maximum in the response to indicator $\varepsilon^{(1)}(t)$, due to the fact that the velocity exceeds the range of indicator $\varepsilon^{(0)}(t)$ frequently, but this is not the case.
One possible explanation for the absence of a visible acceleration phase in the response to indicator $\varepsilon^{(1)}(t)$ at Breitscheid is the frequency at which the sample velocity distribution occurs, i.e. the latter is a rare case on average and the velocity does not exceed the range of indicator $\varepsilon^{(1)}(t)$ as often as in the sample in Fig. 6(a). Due to the fact that the response functions detect an average change, another possible explanation would be a difference in frequency of the traffic oscillation itself. Therefore, we take a closer look at the data sample in Fig. 7. Counting the number of traffic oscillations, i.e the transition from one velocity minimum into the next velocity minimum (which is around 10 km/h for Breitscheid and around 25 km/h for A3 in Fig. 7), during the first 30 minutes of the rush hours one finds five oscillations in case of motorway A3 and three oscillations in case of interchange Breitscheid. Based on the sample, this would imply that the temporal frequency of the traffic oscillation on motorway A3 is almost twice as high as the one at interchange Breitscheid.
@@
Apart from the above statements, which could already be derived from the response functions, one may be able to establish a connection between the width of the valleys in the course of the responses and the periodicity of the traffic oscillation. Since the valleys (hills) represent a sequence of deceleration and acceleration (acceleration and deceleration) one may expect their width (regarding the $\tau$-axis) to coincide with the average time it takes to traverse such a sequence during congestion. To test this assumption we again consider the data sample in Fig. 7 and compare the period length of the traffic oscillation with the width of the valleys in the course of the responses. Therefore, we take into account the range of velocities depicted in the samples. For interchange Breitscheid we find a period of roughly 10 minutes for the traffic oscillation with occurring velocities between 5 km/h and 35 km/h in Fig. 7(a). Comparing this with the corresponding indicators $\varepsilon^{(0)}(t)$ and $\varepsilon^{(1)}(t)$ in Fig. 4(a) we find the width of the valleys is also roughly 10 minutes and therefore in accordance with the period. With the exception of the density response, whose signal curve is blurred, the coincidence is also found for the flow. In case of motorway A3 the sample in Fig. 7(b) mostly shows velocities between 25 km/h and 60 km/h and an oscillation period about 5 minutes. A comparison of the latter with the width of the valleys in the responses to indicators $\varepsilon^{(1)}(t)$ and $\varepsilon^{(2)}(t)$ in Fig. 4(c) also shows the coincidence in case of velocity, flow and density. These findings corroborate the above results regarding the temporal frequency of the traffic oscillations.
@@
Lastly, we would like to make a brief remark on the classification of the found congestion patterns according to three-phase traffic theory. A full analysis on the identification of traffic phases or phase transitions is out of the scope of this study, but it is worth showing a small glimpse of the potential that response functions offer for this purpose. The above findings suggest that the congestion at interchange Breitscheid and on motorway A3 are very likely to be wide moving jams. This conclusion can be traced back to two fundamental properties of wide moving jams [14]. On the one hand a wide moving jam propagates upstream through any traffic state and bottleneck with a constant mean velocity of its downstream jam front. On the other hand wide moving jams have a significant spatial (longitudinal) expansion. In case of interchange Breitscheid the responses indicate congestion with a spatial expansion of about $4.2$km (distance between sections 1 and 5) which propagates through a bottleneck at section 6 (see Sec. 4.2 and Fig. A1 in Appendix A.1). The responses to sections on motorway A3 indicate that congestion propagates an even longer distance, here the spatial extension is about $8.8$km (distance between sections 14 and 19). These findings can be derived from the calculated responses without taking into account any additional data and deliver strong evidence for the classification of the congestion as a wide moving jam.
@@ 4.4 The propagation velocity of congestion
We aim to derive the propagation velocity of congestion from the responses to further strengthen the conclusions from Sec. 4.3. Figure 8 depicts the velocity responses for sections on motorway A3 already shown in Fig. 4(c) broken down by the individual indicators. This representation emphasizes the propagation nature of the minima and maxima, and therefore the propagation of congestion, in a very clear fashion. This clearly striking representation by itself gives rise to the question if it is possible to derive the propagation velocity $v_{c}$ of congestion from our results, which values typically lie between 15 km/h and 20 km/h [14, 27]. To answer the question we determine the propagation velocity of congestion from all types of responses for interchange Breitscheid and motorway A3. A complete overview of all responses for both locations in the representation of Fig. 8 is given in Fig. A2 in the Appendix A.2. To determine the propagation velocities, we work out the positions of the corresponding local minima (or maxima) in the response curves for two sections $k$ and $l$ to retrieve the time lag difference $\Delta\tau^{(\sigma)}_{kl}$. Dividing the distance $d_{kl}$ between the two sections by the corresponding time lag difference determines the propagation velocity $v_{\text{prop}}^{(x)}$ of the congestion:
@@
Figure 9 shows the propagation velocities for both considered cases. Regarding the result of interchange Breitscheid we point out that section 6 is a special case due to its location (see Secs. 3 and 4.2, and Fig. A1 in Appendix A.1), and therefore separated with a dashed line in Fig. 9 (a). In general, we find the values of the propagation velocities either very close to or within the typical interval of $v_{c}$. In the majority of cases the values of $v_{\text{prop}}$ decrease into the numerical range of $v_{c}$ with increasing distance to the indicator section $j$. In other cases, the values of $v_{\text{prop}}$ derive from this behavior by scattering, e.g. $v_{\text{prop}}^{(\rho)}$ for Breitscheid in Fig. 9(a) (last row) as well as $v_{\text{prop}}^{(q)}$ for motorway A3 in Fig. 9(b) (second row). A comparison of these cases with the corresponding responses in Fig. 4 (or Fig. A2) shows that the scattering can be ascribed to the quality of the responses. The shape of the valleys and position of the extrema are not as clearly manifested as in other cases, therefore it is difficult to accurately calculate the related propagation velocities. Nevertheless, the numerical values lie within the typical range of $v_{c}$. Unfortunately, it is not possible to derive a generic behavior on the basis of the results in Fig. 9 due to the scattering of values.
@@
In view of the fact that we derive the propagation velocities from the averaged response functions in an initial approach, we consider the results to be satisfying. Even though the velocities were determined for all indicators separately, the results are close to each other. The results are not conclusive enough to derive a generic behavior regarding the dependence of the distance to the indicator section $j$. However, it is not our goal to carry out a complete analysis within the scope of this contribution, but the above results demonstrate a proof of concept. In case of the classification of the congestion as wide moving jams, the results do not argue against it.

### v2

@@ 4.2 On- and off-ramp propagation of congestion and generic response features
Besides the results of Sec. 4.1 we are interested in the capability of response functions to detect the propagation of congestion through on- and off-ramps into other parts of the motorway network. Therefore we take a look at two cases at interchanges Breitscheid and Herne. In both cases we calculate the responses of spatially subsequent sections along an on-ramp or off-ramp, respectively, to an indicator section on a congested motorway. For interchange Breitscheid we chose section 1 as the indicator section and sections 2 and 6-8 as responding sections, where section 7 is located on the on-ramp leading traffic from motorway A3 onto motorway A52 and section 8 located on motorway A3 (see Fig. 2(a)). In case of interchange Herne we select section 9 as the indicator section and sections 10, 12 and 13 as responding sections, where section 12 is located on the on-ramp leading traffic from motorway A43 onto A42 (see Fig. 2(b)). Figure 5 contains the resulting responses for both cases.
As expected, all types of responses are identical for section 2 at interchange Breitscheid in
@@
Fig. 5(a) (compared with Fig. 4(a)). Furthermore, congestion can be seen moving through the on- or off-ramp, as seen in the responses of all three traffic variables at sections 6 and 7. In this case, the indicating minima (maxima) for $\varepsilon^{(0)}(t)$ are the most distinct. The responses for section 8 do not show any local minima or maxima, indicating that the congestion does not propagate to section 8. It is also noteworthy that there is no available detector between sections 7 and 8.
@@
In the case of the Herne interchange, a distinct indication of congestion can be observed propagating through the on- or off-ramp at section 12, as borne out by the velocity responses to indicator $\varepsilon^{(0)}(t)$ in Fig. 5(b). A comparison between the latter and the corresponding flow responses demonstrates that the flow data is unable to identify the propagation of congestion in this case. The flow responses of sections 12 and 13 exhibit a highly fluctuating course without any local extrema. In contrast, the responses of section 10 exhibit local minima (and maxima in the case of indicator $\varepsilon^{(2)}(t)$). An anomaly is observed when considering the corresponding density responses. The density responses of sections 12 and 13 to indicator $\varepsilon^{(0)}(t)$ exhibit local maxima that align with the courses of the corresponding velocity responses, indicating the propagation of congestion. In contrast, the density response at section 10 exhibits a local minimum, contrary to the expected local maximum.
@@
Notwithstanding the possibility to clearly identify congestion propagation through on- or off-ramps in both cases in Fig. 5, we find the unexpected result of the density response exhibiting a maximum for section 10 at interchange Herne. Lacking any obvious traffic-related reason or systematic error, the unexpected result may be caused by missing values in the analyzed data. For the indicator section 9 about 14% of that data are missing. For section 10 only 4% of the flow data and 9% of the velocity data are missing. The fluctuating courses of the flow responses also suggest a possible impact of low data quality on the anomalous results. Unfortunately, we cannot corroborate this point due to a lack of other available data, e.g. on construction sites. At this point, it is clear that data quality and detector availability limit our study. Both the characteristics of the minima (and maxima) and the fluctuations in the response functions depend directly on the data quality. In addition, a higher resolution of the data could provide more accurate results, but the measurements of the detectors take place over a time interval of one minute. For this reason, the detectors themselves limit our study.
@@
Apart from the congestion propagation, the responses in Fig. 5 and Fig. 4 show similar courses for large time lags $\tau$. Comparing all the cases studied, we find that the velocity responses increase while the flow and density responses decrease with increasing time lag $\tau$. Our results suggest that this behavior is a generic feature for the individual response types induced by the indicators defined in Eq. (6). These generic courses can be explained by the usual traffic situation represented by the averaged data in Fig. 3. With increasing time, the traffic situation during rush hours gradually calms down, i.e. the number of traffic participants decreases. As a consequence, a decrease in density and flow allows higher velocities for the remaining traffic participants until the end of the rush hours. The generic feature we found is suitable for all responses in our study.
@@ 4.3 Differences in congestion patterns
In Sec. 4.1 we observe a deviation in the responses to indicator $\varepsilon^{(2)}(t)$ in comparison to the other indicators. The aforementioned deviant behavior clearly indicates an acceleration phase of the vehicles to moderate velocities for sections at interchange Herne and motorway A3 (in Fig. 4 (b) and (c)), but not in the case of interchange Breitscheid (in Fig. 4 (a)). Why is this so? From the perspective of traffic dynamics, the results indicate the presence of different congestion patterns. In the case of the A3 motorway or the interchange at Herne, vehicles are able to accelerate to higher velocities within the range of indicator $\varepsilon^{(2)}(t)$ before decelerating again due to the congestion. Consequently, larger gaps between traffic participants must be present, otherwise vehicles would be unable to accelerate to a higher velocity range. Therefore, the simultaneous occurrence of the extrema in the responses to indicators $\varepsilon^{(0)}(t)$ and  $\varepsilon^{(2)}(t)$ suggests that the congestion pattern at these locations is characterized by a stop-and-go nature, exhibiting acceleration-deceleration cycles. In contrast, the responses to indicators $\varepsilon^{(1)}(t)$ and $\varepsilon^{(2)}(t)$ at interchange Breitscheid do not exhibit any indication of an acceleration phase, i.e. local maxima, during the critical congestion period. This finding suggests that the congestion pattern is more dense and synchronized, with vehicles moving at low velocities (in the range of $\varepsilon^{(0)}(t)$ and $\varepsilon^{(1)}(t)$) while maintaining small distances to vehicles ahead of them.
@@
This does not necessarily rule out a stop-and-go pattern of congestion, but the effect is not strong enough to be captured by the responses.
Given the nature of the data at our disposal, it is not possible to verify the aforementioned assumptions about the congestion pattern with information regarding the distances between traffic participants. Consequently, we consider the raw data. Figure 6 illustrates a sample of the daily data, according to Eq. (4), for all locations under consideration. In all cases, the time period of the traffic breakdown can be discerned by examining the trends in velocity and density. Due to its fluctuating nature, the impact on traffic flow is not as readily apparent as with other observables. What the average responses already indicate is revealed by the velocity data (first column in Fig. 6). While the velocity at section 3 at interchange Breitscheid is nearly always below the range of indicator $\varepsilon^{(2)}(t)$, the velocities at sections 12 and 15 at interchange Herne and on motorway A3, respectively, frequently reach (and exceed) the range of indicator $\varepsilon^{(2)}(t)$. Although a traffic oscillation, or stop-and-go pattern, is evident in the velocities at all three locations, the corresponding responses in Fig. 4 only show an acceleration phase, or local maxima, for indicator $\varepsilon^{(2)}(t)$ in the case of interchange Herne and motorway A3. Considering the velocity data for section 3 at Breitscheid (in Fig. 6), one is led to anticipate a local maximum in the response to indicator $\varepsilon^{(1)}(t)$ due to the fact that the velocity frequently exceeds the range of indicator $\varepsilon^{(0)}(t)$. However, this is not the case.
One possible explanation for the absence of a visible acceleration phase in the response to indicator $\varepsilon^{(1)}(t)$ in Breitscheid is the frequency with which the sample velocity distribution occurs, i.e. the latter is on average a rare case and the velocity does not exceed the range of indicator $\varepsilon^{(1)}(t)$ as often as in the example in Fig. 6(a). Another possible explanation is a difference in the frequency of the traffic oscillation itself. Therefore, we take a closer look at the data sample in Fig. 7. By enumerating the number of traffic oscillations, or the transition from one velocity minimum to the next (which, as illustrated in Fig. 7, is approximately 10 km/h for Breitscheid and approximately 25 km/h for the A3), it is observed that there are five oscillations for the A3 and three oscillations for the Breitscheid interchange during the initial 30 minutes of the rush hour. Therefore, the temporal frequency of the traffic oscillations on the A3 is approximately twice that observed at the Breitscheid interchange.
@@
Furthermore, a possible correlation may be established between the full width of the minima in the course of the responses and the periodicity of the traffic oscillation, in addition to the aforementioned statements. Given that the valleys or minima (hills or maxima) represent a sequence of deceleration and acceleration (acceleration and deceleration), it may be expected that their width (in relation to the $\tau$-axis) corresponds to the average time taken to traverse such a sequence during congestion. To test this assumption, we again consider the data sample shown in Fig. 7 and compare the period length of the traffic oscillation with the width of the valleys in the course of the responses. Accordingly, the range of velocities depicted in the samples is taken into account. With regard to the interchange at Breitscheid, a period of approximately ten minutes is observed for the traffic oscillation, with velocities ranging from 5 to 35 km/h, as illustrated in Fig. 7(a). A comparison of this with the corresponding indicators $\varepsilon^{(0)}(t)$ and $\varepsilon^{(1)}(t)$ in Fig. 4(a) reveals that the width of the valleys is also approximately 10 minutes, thus confirming the period. With the exception of the density response, for which the signal curve is blurred, the aforementioned coincidence is also observed in the flow. With regard to the A3 motorway, the sample depicted in Fig. 7(b) predominantly exhibits velocities between 25 and 60 km/h, with an oscillation period of approximately five minutes. A comparison of the latter with the width of the valleys in the responses to indicators $\varepsilon^{(0)}(t)$ and $\varepsilon^{(1)}(t)$ in Fig. 4(c) also demonstrates the coincidence in the case of velocity, flow, and density. These findings substantiate the aforementioned results for the temporal frequency of the traffic oscillations.
@@
In conclusion, we briefly comment on the classification of the identified congestion patterns according to the three-phase traffic theory. A comprehensive examination of the identification of traffic phases or phase transitions is beyond the scope of this study. However, it is worthwhile to hint at the potential offered by response functions for this purpose. The aforementioned findings indicate that the congestion observed at interchange Breitscheid and on the A3 motorway is likely to be a wide-moving jam. This conclusion can be traced back to two fundamental properties of wide moving jams [13]. On the one hand, a wide moving jam propagates upstream through any traffic state and bottleneck with a constant mean velocity of its downstream jam front. On the other hand, wide moving jams exhibit a notable longitudinal expansion. In the case of the interchange at Breitscheid, the responses indicate congestion with a spatial expansion of approximately $4.2$ km (the distance between sections 1 and 5), which propagates through a bottleneck at section 6 (see Sec. 4.2 and Fig. A1 in Appendix A.1). The responses to sections on the A3 motorway indicate that congestion propagates over an even longer distance, with a spatial extension of approximately $8.8$ km (between sections 14 and 19). These findings follow from the calculated responses without additional data, providing compelling evidence for the classification of the congestion as a wide moving jam.
With regard to the differentiation of congestion patterns, our study is not only limited by the quality of the data. Unfortunately the relationship between the widths of the local minima and the traffic oscillation cannot be established using the averaged data in Fig. 3. Furthermore, the relationship is clearer for detectors close to the reference detector than for detectors that are further away. The width of the minima increases slightly for more distant detectors. Therefore, our approach is locally limited.
@@ 4.4 The propagation velocity of congestion
Figure 8 depicts the velocity responses for sections on motorway A3, as previously illustrated in Fig. 4(c), broken down by the individual indicators. This representation serves to emphasise the propagation nature of the minima and maxima, and therefore the propagation of congestion, in a particularly clear manner. This striking representation prompts the question of whether the propagation velocity $v_{c}$ of congestion can be derived from our findings, which typically ranges between 15 km/h and 20 km/h [13, 26]. To answer the question, we approach to determine the propagation velocity of congestion from all types of responses for the Breitscheid interchange and the A3 motorway. A complete overview of all responses for both locations in the representation of Fig. 8 is given in Fig. A2 in the Appendix A.2. To determine the propagation velocities, we work out the positions of the corresponding local minima (or maxima) in the response curves for two sections $k$ and $l$ to obtain the time lag difference $\Delta\tau^{(\sigma)}_{kl}$. Dividing the distance $d_{kl}$ between the two sections by the corresponding time lag difference determines the propagation velocity $v_{\text{prop}}^{(x)}$ of the congestion:
@@
Figure 9 shows the propagation velocities for the two cases considered. Regarding the result of the Breitscheid interchange, we note that section 6 is a special case due to its location (see Secs. 3 and 4.2, and Fig. A1 in Appendix A.1), and is therefore separated by a dashed line in Fig. 9 (a). In general, we find the values of the propagation velocities either very close to or within the typical interval of $v_{c}$. In the majority of cases the values of $v_{\text{prop}}$ decrease into the numerical range of $v_{c}$ with increasing distance to the indicator section $j$. In other cases, the values of $v_{\text{prop}}$ derive from this behavior by scattering, e.g. $v_{\text{prop}}^{(\rho)}$ for Breitscheid in Fig. 9(a) (last row) as well as $v_{\text{prop}}^{(q)}$ for motorway A3 in Fig. 9(b) (second row). A comparison of these cases with the corresponding responses in Fig. 4 (or Fig. A2) shows that the scattering can be ascribed to the characteristics of the responses. The shape and position of the local minima are not as clearly manifested as in other cases, making it difficult to accurately calculate the associated propagation velocities. Nevertheless, the numerical values are consistent with the typical range of $v_{c}$. Unfortunately, it is not possible to derive a generic behavior on the basis of the results in Fig. 9.
@@
Considering that we derive the propagation velocities from the average response functions in a first approach, the results are meaningful. Although the velocities were determined separately for all indicators, the results are close to each other. While these results are not conclusive enough to derive a general behaviour regarding the dependence of the distance on the indicator section $j$, the results provide a proof of concept. In the case of the classification of the congestion as a wide moving jam, the results do not argue against it.